After retirement, Nguyen wants to open a Corona Spa. He is planning to retire in the next 30

years. He wants to be able to draw out of his savings 30 years after retirement an amount of

300 mil VNĐ per year. How much would he has to pay in equal annual deposits until

retirement to meet his objectives? Assume interest remains at 3%


1
Expert's answer
2022-01-12T08:33:55-0500

FV=300

r=0.03

t=30

"FV=A\frac{(1+r)^n-1}{r}"


"300=A\frac{(1+0.03)^{30}-1}{0.03}=6.3"